OpenStudy (anonymous):

What is the variance of the following data? If necessary, round your answer to two decimal places. 6, 7, 10, 11, 11, 13, 16, 18, 25 31.11 33.89 32.56 34.9 I'm stuck between 31.11 and 34.9 I think it's A am I right?

OpenStudy (anonymous):

It would be A right?

OpenStudy (tkhunny):

Are the data an entire population or a sample?

OpenStudy (anonymous):

I have no idea

OpenStudy (tkhunny):

There you go. Without that information, you are stuck. It's usually wiser to select the greater of the two if there is no evidence sufficient to decide.

OpenStudy (anonymous):

So it's D?

OpenStudy (anonymous):

Because when it's done I either get 35 or 31.11, 35 isn't an option, so I feel like it's A

OpenStudy (anonymous):

Fisrt find mean value: M = (6+7+10+11+11+13+16+18+25)/9 M = 13 Now find the sum of square deviations: D = (6-13)^2+(7-13)^2+(10-13)^2+(11-13)^2+(11-13)^2+(13-13)^2+(16-13)^2+(18-13)^2+(25-13)^2 D = 280 Now divide D: D/9 280/9 = 31

OpenStudy (anonymous):

it's A

OpenStudy (anonymous):

280/9 = 31.11

OpenStudy (anonymous):

So just follow this algorithm: 1) Find the mean value of your data; 2) Find all deviations, i.e. take each data in your table a subtract the mean on it; 3) Sum the squares of each subtration on the step 2; 4) Divide the result of step 3 by the number of data in your table; Done, this resutlt is the variance of your data set

OpenStudy (anonymous):

Oh yay! Thanks!

OpenStudy (anonymous):

;)